# Characterizing variability in complex network community structure with a recursive significance clustering scheme

### v1

@@ V Conclusions
We introduced a recursive significance clustering scheme to extract robust cores from a network, defined as subsets of nodes with stable joint community assignments under network perturbation. Cores are the fundamental building blocks of community structure, delineating structure variability. This method also inherently distinguishes the nodes with noisy community assignments, allowing them to be filtered to facilitate meaningful comparisons of network partitions. The coalescence frequency of a set of cores, which we define as the proportion of ensemble solutions in which only those cores unite to form a community, illuminates the structural patterns present in the solution landscape, thereby indicating network features well-supported by the data.
@@
Studies using community detection to guide a scientific pursuit will benefit from a cognizance of cores to ensure analyses are not conducted on only one network partition, which may exhibit dissimilar structure from other partitions of similar merit, and of only one network sample. We readily note that while our recursive significance clustering technique has been presented here in the context of complex network community structure, it is freely extensible for an examination of any clustered data.
VI Materials and methods
VI.1 Empirical network
The described AK–AK flight network was constructed from the 2017, 2020, and 2023 T-100 Domestic Segment Data from the Bureau of Transportation Statistics. Entries are flight segments, defined as a pair of domestic US airports served by a single stage of at least one flight, with or without passengers during the reporting year. Available data for each segment includes its departure and arrival aerodromes and states, and the total annual flight count. For a given year, we consider only the airports in AK with at least one flight to or from another airport in AK. We use IATA, FAA, and ICAO identifiers to obtain airport GPS coordinates.
@@ VI.1 Empirical network
We mathematically formulate each temporal snapshot of the network (graph) as $G=\left(V,E,W\right)$, where $V$ is the network’s set of nodes (vertices), $E\subseteq V\times V$ is its set of directed links (edges), and $W\colon E\to\mathbb{R}^{+}$ is a weighting function that assigns positive, real weights to links. Nodes in the AK–AK flight network represent aerodromes, and directed edges flight segments, weighted by annual flight count.
VI.2 Network perturbation
@@
We aim now to somehow perturb the atomic components of $G$ repeatedly to generate a collection of $N=1000$ replicate networks $\mathcal{G}=\{G^{*}_{1},\ldots,G^{*}_{N}\}$. For many complex networks, it makes little sense to discuss the network having a different set of nodes. Zachary’s Karate Club, for example, is an idiosyncratic network that would be unrecognizable without the instructor, or with an additional officer; the same is true for our flight network. So as to not undermine the individual characteristics of nodes, we turn then to modifying the depicted relationships between them, i.e., link weights [32].
@@
Convenient for networks that preclude replicate observations, we make use the statistical bootstrap [42] to produce replicate networks by parameterizing and resampling link weights [32]. Treating flights as independent events, we resample each link’s weight from a Poisson distribution with a rate parameter equal to that link’s weight in the empirical network. In this manner, each replicate network $G^{*}_{i}=\left(V,E,W^{*}_{i}\right)$ varies by weighting function where the weight of each link $e\in E$ is iid such that $W^{*}_{i}(e)\sim\mathrm{Pois}(W(e))$.
@@
We stress that a plethora of other perturbation/resampling methods are acceptable. For example, introducing additive noise over link weights has the advantage of being directly interpreted as uncertainty in the empirical network configuration in many contexts [25]. Links may be repositioned [27], or a random sampling of them may be taken [43], or other suitable methods to capture variability of the empirical network. Indeed, if many replicate observations of the network were available, these could be used in lieu of generating artificial networks.
@@
In the introductory example of Zachary’s Karate Club (fig. 1), instead of perturbing the network to generate a partition ensemble, we effectively perturbed the stochastic community detection algorithm operating on a static network through the use of different random seeds. While beyond the scope of this work, we acknowledge the potential utility of this method in identifying structure common among different strategies of community detection.
VI.3 Community detection
@@
We now require a community description $M^{*}_{i}$ that is a partition of the node set $V$ for each network realization $G^{*}_{i}\in\mathcal{G}$ to generate an ensemble of partitions =$\mathcal{M}=\{M^{*}_{1},\ldots,M^{*}_{N}\}$. We consider non-overlapping as opposed to hierarchical community structure such that each node is assigned to one and only one community in each community description for ease of interpretation of results.
@@
Leveraging that our flight network is both directed and weighted, and can be analogized as a model of air traffic flow, we use the stochastic search algorithm Infomap [12] making use of the map equation framework [7] to detect community structure. The map equation quantifies the quality of a network partition by its ability to compress information flow on the network as a means of identifying regularities in network structure with respect to flow [15]. Communities under the map equation regime then correspond to structures of persistent flow, implying strong internal mixing and coherence [44]. To tune detected communities to a desirable scale for analysis, we set a variable [45] Markov time [46, 47] of 3. Other community detection algorithms more befitting the empirical network under study may be used [6].
VI.4 Core search
@@
We wish to find the cores of the network $G$ with significance level $\alpha=0.05$, i.e., subsets of nodes that are jointly assigned to the same community over a $(1-\alpha)$-fraction of community descriptions in the solution ensemble $\mathcal{M}$. This is done in a recursive manner, such that the largest core of the network is identified first, then the next largest, and so on until all desired cores are delineated. To define what is meant by the ”largest” core, we must specify a measure of size of a subset of nodes $\mu:2^{V}\to\mathbb{R}^{+}$, which we take simply to be the cardinality of the set. This choice, in essence, treats each node as having equal importance in the delineation of cores.
@@
Following the seminal work on significance clustering of [32], the score of an arbitrary candidate core $C\subseteq V$ is denoted $s(C)$ and is the size of the core $\mu(C)$ minus a penalty to ensure its nodes cluster together over a $(1-\alpha)$-fraction of ensemble solutions. To calculate the penalty, for every community description $M^{\prime}_{i}\in\mathcal{M}$, we find the community $m\in M^{\prime}_{i}$ that most closely matches the candidate core $C$. The mismatch nodes $C\setminus m$ that are assigned to $C$ but not to $m$ break the constraint that the core must cluster together. The size of this mismatch node set $\mu(C\setminus m)$ in every community description is summed, excepting the $\alpha$-fraction producing the largest penalties to relax the requirement. This quantity is the penalty associated with $C$.
@@
To approximate the largest core, we employ simulated annealing [48]. To begin, on sweep $t=0$ with temperature $T_{0}=10$, a candidate core $C\subseteq V$ is randomly initialized. A node $v\in V$ is drawn at random, and has its membership to the candidate core $C$ flipped to create a trial core $C^{\prime}$. Per the Metropolis–Hastings algorithm [49, 50], this trial core is accepted as the candidate core $C\leftarrow C^{\prime}$ with probability equal to $\min\left\{1,\,\exp((s(C^{\prime})-s(C))/T_{t})\right\}$.
@@
The generation of candidate cores is repeated with replacement of drawn nodes as many times as there are nodes to draw from to complete one full sweep. After a full sweep, the temperature is cooled on an exponential schedule $T_{t}=T_{0}r^{t}$, where $r\in\left(0,1\right)$ is a cooling rate parameter we set ad hoc to $r=0.9$. Cooling continues until no trial core is accepted on a sweep, at which point the candidate core $C$ is taken as an approximation to the largest core.
@@ Appendix A Temporal snapshots
Figure 5: Cores of the 2020 temporal snapshot of the AK–AK flight network. (a) Spatially-embedded cores. Nodes with unstable community assignments are shaded grey. (b) Core coalescence frequency and stability.
Figure 6: Cores of the 2023 temporal snapshot of the AK–AK flight network. (a) Spatially-embedded cores. Nodes with unstable community assignments are shaded grey. (b) Core coalescence frequency and stability.

### v2

@@ VI Conclusions
Here we have introduced a recursive significance clustering scheme to identify the robust features the variability of community structure in complex networks. We extracted the cores of networks as the subsets of nodes that have stable joint community assignments under network perturbation. Cores serve as atomic building blocks that are arranged in different configurations to form a community structure landscape but are nearly indivisible themselves. Based on these properties of cores, we defined the core coalescence frequency and stability metrics as the frequency that cores unite into communities and nodes into cores, respectively. This method also inherently distinguishes the nodes with noisy community assignments, allowing them to be filtered to facilitate meaningful comparisons of network partitions.
@@
Studies using community detection to guide a scientific pursuit will benefit from a cognizance of cores to ensure analyses are not conducted on only one network partition of only one network sample, and instead embrace uncertainty or noise in the network’s configuration to demarcate the robust features of its structure. We note that while our recursive significance clustering technique has been presented here in the context of complex network community structure, its premise is perfectly suitable for an examination of any clustered data. It would benefit from future development extending its use to applications of hierarchical clustering.
We look forward to a full comparison of network cores from recursive significance clustering and consensus clustering solutions in future work. We foresee the use of cores as building blocks to form a consensus solution of robust community structure according to the needs of one’s application.
VII Materials and methods
VII.1 Network
@@
We mathematically formulate each network as $G=\left(V,E,W\right)$, where $V$ is the network’s set of nodes, $E\subseteq V\times V$ is its set of (directed) links, and $W\colon E\to\mathbb{R}^{+}$ is a weighting function that assigns positive and real weights to links. When the network is unweighted, the weight of each edge is equal to one. When it is undirected, a link $e_{u,v}\in E$ between nodes $u,v\in V$ is equivalent to the link $e_{v,u}$. To simplify notation, some link that is not a component of the network $e_{u,v}\not\in E$ is assigned a weight of zero $W(e_{u,v})=0$.
AK–AK flight network
The described AK–AK flight network was constructed from the 2017, 2020, and 2023 T-100 Domestic Segment Data from the US Bureau of Transportation Statistics. Entries are flight segments, defined as a pair of domestic US airports served by a single stage of at least one flight during the reporting year. For a given year, we consider only the airports in AK with at least one flight to or from another airport in AK. We use IATA, FAA, and ICAO identifiers to obtain airport GPS coordinates. Each temporal snapshot of the AK–AK flight network is a separate network with nodes representing aerodromes and directed edges flight segments weighted by annual flight count.
Synthetic network
We generate LFR benchmark networks [48] with variable mixing parameter $\mu$ and average degree $\langle k\rangle$. With 1000 nodes, we set the maximum degree and community size equal to 100 to practically synthesize the network [50, 51]. We use degree and community size distribution exponents of -3 and -2, respectively, which are typical values found in real networks [48], including the AK–AK flight network. We generate three networks with different random seeds per parameterization and report the mean of results.
VII.2 Network perturbation
@@
We aim now to somehow perturb the atomic components of $G$ repeatedly to generate a collection of $N=1000$ replicate networks $\mathcal{G}=\{G^{*}_{1},\ldots,G^{*}_{N}\}$. For many complex networks, it makes little sense to discuss the network having a different set of nodes. Zachary’s Karate Club, for example, is an idiosyncratic network that would be unrecognizable without the instructor or with an additional officer; the same is true for our flight network. So as to not undermine the individual characteristics of the nodes, we modify the relationships between them, that is, links [36].
AK–AK flight network
@@
Convenient for networks that preclude replicate observations, we make use of the statistical bootstrap [52] to produce replicate networks by parameterizing and resampling link weights [36]. Treating flights as independent events, we resample each link’s weight from a Poisson distribution with a rate parameter equal to that link’s weight in the empirical network. In this manner, each replicate network $G^{*}_{i}=\left(V,E,W^{*}_{i}\right)$ varies by weighting function, where the weight of each link $e\in E$ is i.i.d. such that $W^{*}_{i}(e)\sim\mathrm{Pois}(W(e))$.
Synthetic network
As we examine the stability of community structure as a function of mixing $\mu$, which is concerned with the position of links between and within communities, we perturb each LFR benchmark network by rewiring a small percentage of its links. For $\Sigma$ undirected and unweighted links, we randomly draw $0.05\Sigma$ links to redirect. One endpoint of each drawn edge is replaced with another node at random; if such an edge already exists in the network it is simply removed since the network is unweighted.
@@
We stress that a plethora of other perturbation/resampling methods are acceptable. For example, introducing additive noise over link weights has the advantage of being directly interpreted as uncertainty in the empirical network configuration in many contexts [24]. Links may be repositioned in alternative fashions [26], added [51], or randomly sampled [53], and other suitable methods may be used to model variability in the networked system. Indeed, if many replicate observations of the network were available, these could be used in lieu of generating artificial networks.
@@
In the introductory example of Zachary’s Karate Club (fig. 1), instead of perturbing the network to generate a partition ensemble, we effectively perturbed the stochastic community detection algorithm operating on a static network through the use of different random seeds [54]. Although beyond the scope of this work, we acknowledge the potential utility of this method in identifying common structure among results from different community detection strategies.
VII.3 Community detection
@@
We now require a community description $M^{*}_{i}$ that is a partition of the node set $V$ for each network realization $G^{*}_{i}\in\mathcal{G}$ to generate a set of partitions =$\mathcal{M}=\{M^{*}_{1},\ldots,M^{*}_{N}\}$. We consider non-overlapping as opposed to hierarchical community structure such that each node is assigned to one and only one community in each community description for ease of interpretation of results. The community detection algorithm that best fits the network under study should be used [6].
AK–AK flight network
@@
Leveraging that our flight network is both directed and weighted, and can be analogized as a model of air traffic flow, we use the stochastic search algorithm Infomap [11] making use of the map equation framework [7] to detect community structure. The map equation quantifies the quality of a network partition by its ability to compress information flow on the network as a means of identifying regularities in network structure with respect to flow [14]. Communities under the map equation regime then correspond to structures of persistent flow, implying strong internal mixing and coherence [55]. To tune detected communities to a desirable scale for analysis, we set a variable Markov time [56, 57, 58] of 3.
Synthetic network
The modularity of a partition $M$ of the undirected network $G$ is
$Q(G,M)=\frac{1}{2\Sigma}\sum_{u,v\in V}\left(W(e_{u,v})-\frac{k_{u}k_{v}}{2\Sigma}\right)\delta(m_{u},m_{v}),$
where $\Sigma$ is the sum of link weights in the network, $k_{u}$ is the strength of node $u$, and $\delta$ is the Kronecker delta function such that $\delta(m_{u},m_{v})$ evaluates to unity if nodes $u$ and $v$ are jointly assigned to a community in $M$ and zero otherwise [59]. The Louvain method of community detection finds a partition $M$ that maximizes $Q$ through a greedy stochastic search [9].
VII.4 Core search
@@
We wish to find the cores of the network $G$ with significance level $\alpha=0.05$, i.e., subsets of nodes that are jointly assigned to the same community over a $(1-\alpha)$-fraction of community descriptions in the solution ensemble $\mathcal{M}$. This is done in a recursive manner, such that the largest core of the network is identified first, then the next largest, and so on until all desired cores are delineated. To define what is meant by the ”largest” core, we must specify a measure of size of a subset of nodes $\nu:2^{V}\to\mathbb{R}^{+}$, which we take simply to be the cardinality of the set. This choice, in essence, treats each node as having equal importance in the delineation of cores.
@@
Following the seminal work on significance clustering of [36], the score of an arbitrary candidate core $C\subseteq V$ is denoted $s(C)$ and is the size of the core $\nu(C)$ minus a penalty to ensure its nodes cluster together over a $(1-\alpha)$-fraction of ensemble solutions. To calculate the penalty, for every community description $M^{\prime}_{i}\in\mathcal{M}$, we find the community $m\in M^{\prime}_{i}$ that contains the most nodes of the candidate core $C$. The mismatch nodes $C\setminus m$ that are assigned to $C$ but not to $m$ break the constraint that the core must cluster together. The size of this mismatch node set $\nu(C\setminus m)$ in every community description is summed, excepting the $\alpha$-fraction producing the largest penalties to relax the requirement. This quantity is the penalty associated with $C$.
@@
To approximate the largest core, we employ simulated annealing [60]. To begin, on sweep $t=0$ with temperature $T_{0}=1$, a candidate core $C\subseteq V$ is randomly initialized. A node $v\in V$ is drawn at random, and has its membership to the candidate core $C$ flipped to create a trial core $C^{\prime}$. Per the Metropolis–Hastings algorithm [61, 62], this trial core is accepted as the candidate core $C\leftarrow C^{\prime}$ with probability equal to $\min\left\{1,\,\exp((s(C^{\prime})-s(C))/T_{t})\right\}$.
@@
The generation of candidate cores is repeated with replacement of drawn nodes as many times as there are nodes to draw from to complete one full sweep. After a full sweep, the temperature is cooled on an exponential schedule $T_{t}=T_{0}r^{t}$, where $r\in\left(0,1\right)$ is a cooling rate parameter we set ad hoc to $r=0.99$. Cooling continues until no trial core is accepted on a sweep, at which point the candidate core $C$ is taken as an approximation to the largest core.
@@
The simulated annealing scheme used here is relatively primitive compared to the modern search algorithms used in community detection. Its run-time performance can be improved to scale more efficiently with network size by incorporating techniques from the literature [6], most valuably heuristic candidate generation and restarts.
@@ Appendix A Temporal snapshots
Figure 6: Cores of the 2020 temporal snapshot of the AK–AK flight network. (a) Spatially embedded cores. Nodes with unstable community assignments are shaded grey. (b) Core coalescence frequency and stability.
Figure 7: Cores of the 2023 temporal snapshot of the AK–AK flight network. (a) Spatially embedded cores. Nodes with unstable community assignments are shaded grey. (b) Core coalescence frequency and stability.